Sandra wants to place a utilitarian door on the back entrance that leads to her patio. The solid plastic laminate door will be shaped like a rectangular solid and will measure 72 3/4 inches long and 32 2/3 inches wide; and will be 1 3/8 inch thick. How many cubic inches of plastic laminate material will be needed for the door? Give solution exactly as a mixed number. Include correct units with your answer.
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16946)

length of door =72 3/4 "= 291/4
32 2/3 " wide= 98/3
thickness = 1 3/8 " 11/8
Volume = 291/4*98/3 * 11/8
= 3267 17/25



Answer by Alan3354(69443)

Length = 72 3/4 = 291/4
Width = 32 2/3 = 98/3
Thickness = 1 3/8 = 11/8
Volume = (291/4)*(98/3)*(11/8)
= 313698/96
= 52283/16
= 3267 11/16 cubic inches
--------------------------
Mixed numbers are not very useful.